It is known that IQ scores form a normal distribution with a mu of 100 and a standard deviation of 15. Given this information, what is the probability of randomly selecting an individual with an IQ score less than 120?

a)9.72%

b)95%

c)Unable to answer this question with information given.

d)90.15%

Answer by Theo(13342)

first you find the z-score.

the z-score is equal to (x-m)/s

x is the x-score
x is the raw score.
m is the mean.
s is the standard deviation.

the z-score in this case is z = (120-100)/15 = 20/15 = 4/3 = 1.33.....

use a z-score calculator to the highest degree of accuracy that it possesses and the probability becomes .9087887181 which is equal to roughly 90.88%.

closest selection you have is 90.15%, even though it's not right on.

a normal distribution is assumed.